OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2

OpenStudy (aravindg):

USE slope-point form \[\large \boxed{y-y_1=m(x-x_1)}\] where m is slope and point is \((x_1,y_1)\)

OpenStudy (anonymous):

Ik i tried but i still cant get the right answer. Ik im wrong becasue its mulitpule choice.

OpenStudy (anonymous):

A. y – 9 = 2(x – 5) B. y – 5 = 2(x + 9) C. y + 9 = 2(x – 5) D. y – 9 = 2(x + 5)

OpenStudy (aravindg):

can you show ur working?